Direct stresses of 120 N/mm^2 (tension) and 90 N/mm^2 (compression) are applied at a particular point in an elastic material on two mutually perpendicular planes the principal



stress in the material is limited to 100 N/mm^2(tension) calculate the allowable value of shear stress at the point on the given planes determine also the value of the other principal stress and the maximum value of shear stress at the point using Mohr's circle




1
Expert's answer
2022-02-17T01:48:02-0500

"\sigma_x=120"

"\sigma_y=-90"

"\sigma_1=100"

Let allowable value of shear stress be denoted "\tau_{xy}"

"\sigma_1=\frac{1}{2}[(\sigma_x+\sigma_y)+\sqrt{(\sigma_x-\sigma_y)^2+4\tau_{xy}^2}"

"100=\frac{1}{2}[(120-90)+\sqrt{(120+90)^2+4\tau_{xy}^2}"

"100=120+2\tau_{xy}"

"\tau_{xy}=\frac{-20}{2}"

=-10

=10N/mm2 compressive force.